A flask with a volume of 272 ml is filled with an unknown gas mixture. When filled, it is found that this gas mixture weights 0.211 g at 785 torr and 27C. What is the average molecular weight of the gas mixture?
Answer #1

Given:
P = 785.0 torr
= (785.0/760) atm
= 1.0329 atm
V = 272.0 mL
= (272.0/1000) L
= 0.272 L
T = 27.0 oC
= (27.0+273) K
= 300 K

find number of moles using:
P * V = n*R*T
1.0329 atm * 0.272 L = n * 0.08206 atm.L/mol.K * 300 K
n = 1.141*10^-2 mol

mass(solute)= 0.211 g


use:
number of mol = mass / molar mass
1.141*10^-2 mol = (0.211 g)/molar mass
molar mass = 18.49 g/mol
Answer: 18.5 g/mol